Number of ions present in 4.25 moles of fe
1
Expert's answer
2020-11-09T14:11:53-0500

A mole of Fe atoms contains 6.02×1023 Fe atoms.

4.25 moles of Fe atoms contains "4.25 \times 6.02 \times 10^{23} = 2.55 \times 10^{24}" atoms